# Sufficient and Necessary Barrier-like Conditions for Safety and Reach-avoid Verification of Stochastic Discrete-time Systems

### v2

@@ IV-A Reach-avoid Verification I
$\begin{split}&h_{i}(\bm{x})=\mathbb{E}_{\pi}[\prod_{j=1}^{i}1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(j))v(\bm{\phi}_{\pi}^{\bm{x}}(i+1))]\end{split}.$
Since $\mathbb{P}_{\pi}(\forall k\in\mathbb{N}.\bm{\phi}_{\pi}^{\bm{x}}(k)\in\mathcal{X}\setminus\mathcal{X}_{r})=0$ for $\bm{x}\in\mathcal{X}\setminus\mathcal{X}_{r}$ and $v(\cdot):\mathbb{R}^{n}\rightarrow\mathbb{R}$ is bounded over $\mathbb{R}^{n}$, we have $\lim_{i\rightarrow\infty}h_{i}(\bm{x})=0$ for $\bm{x}\in\mathbb{R}^{n}$ and consequently, $v(\bm{x})=V(\bm{x})$ over $\mathbb{R}^{n}$. ∎
@@
Since $v(\bm{x})$ satisfies (10), we can obtain, following the induction of (9) by replacing “=” with $\leq$”, that
$\begin{split}v(\bm{x})\leq V(\bm{x})+1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{x})\lim_{i\rightarrow\infty}h_{i}(\bm{x})\end{split}$
for $\bm{x}\in\mathbb{R}^{n}$, where $V(\cdot):\mathbb{R}^{n}\rightarrow\mathbb{R}$ is the value function defined in (6). Since $v(\bm{\phi}_{\pi}^{\bm{x}}(i+1))\leq 0$ when $\bm{\phi}_{\pi}^{\bm{x}}(i+1)\in\mathbb{R}^{n}\setminus\mathcal{X}$, we have
$\begin{split}h_{i}(\bm{x})=\mathbb{E}_{\pi}[\prod_{j=1}^{i}1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(j))v(\bm{\phi}_{\pi}^{\bm{x}}(i+1))]\leq\mathbb{E}_{\pi}[\prod_{j=1}^{i}1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(j))w_{i+1}(\bm{x})],\end{split}$
where $w_{i+1}(\bm{x})=1_{\mathcal{X}}(\bm{\phi}_{\pi}^{\bm{x}}(i+1))v(\bm{\phi}_{\pi}^{\bm{x}}(i+1))$. Also, since $v(\cdot):\mathbb{R}^{n}\rightarrow\mathbb{R}$ is bounded over $\mathcal{X}$ and $\mathbb{P}_{\pi}(\forall k\in\mathbb{N}.\bm{\phi}_{\pi}^{\bm{x}}(k)\in\mathcal{X}\setminus\mathcal{X}_{r})=0$ for $\bm{x}\in\mathcal{X}\setminus\mathcal{X}_{r}$, we conclude $\lim_{i\rightarrow\infty}h_{i}(\bm{x})=0$ for $\bm{x}\in\mathbb{R}^{n}$. Consequently, $v(\bm{x})\leq V(\bm{x})$ for $\bm{x}\in\mathbb{R}^{n}$.
@@
Without Assumption 1, we cannot use condition (10) to justify lower bounds in the reach-avoid verification, since we cannot guarantee $\lim_{i\rightarrow\infty}h_{i}(\bm{x})=0$ for $\bm{x}\in\mathcal{X}\setminus\mathcal{X}_{r}$.
@@ IV-B Reach-avoid Verification II
The subsection will formulate a necessary and sufficient barrier-like condition for the reach-avoid verification without Assumption 1. Instead, another assumption that $\epsilon_{2}$ is strictly smaller than the exact reach-avoid probability $\mathbb{P}_{\pi}(RA_{\bm{x}_{0}})$ is imposed. Like the one in Subsection IV-A, this condition is constructed via relaxing a Bellman equation, which is derived from a discounted value function.
@@
$\begin{split}\sum_{k=m+1}^{M}\mathbb{E}_{\pi}[\gamma^{k}\prod_{j=0}^{k-1}1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(j))1_{\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(k))]\leq\sum_{k=m+1}^{M}\mathbb{E}_{\pi}[\prod_{j=0}^{k-1}1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(j))1_{\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(k))]\end{split}$
@@
Since $v(\bm{x})$ satisfies the Bellman equation (12), we have
$\begin{split}&v(\bm{x})=1_{\mathcal{X}_{r}}(\bm{x})+\gamma 1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{x})\mathbb{E}_{\bm{\theta}}[v(\bm{f}(\bm{x},\bm{\theta}))]\\ &=1_{\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(0))+\gamma 1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(0))\mathbb{E}_{\bm{\theta}}[v(\bm{\phi}_{\pi}^{\bm{x}}(1))]\\ &=1_{\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(0))+\gamma 1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(0))\mathbb{E}_{\pi}[1_{\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(1))+\gamma 1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(1))\mathbb{E}_{\bm{\theta}}[v(\bm{f}(\bm{\phi}_{\pi}^{\bm{x}}(1),\bm{\theta}))]]\\ &=\mathbb{E}_{\pi}[1_{\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(0))+\gamma 1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(0))1_{\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(1))]+\gamma^{2}1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(0))\mathbb{E}_{\pi}[1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(1))v(\bm{\phi}_{\pi}^{\bm{x}}(2))]\\ &=\cdots\\ &=\mathbb{E}_{\pi}[1_{\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(0))+\sum_{i\in\mathbb{N}_{\geq 1}}\gamma^{i}\prod_{j=0}^{i-1}1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(j))1_{\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(i))]+1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{x})\lim_{i\rightarrow\infty}h_{i}(\bm{x})\\ &=\tilde{V}_{\gamma}(\bm{x})+1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{x})\lim_{i\rightarrow\infty}h_{i}(\bm{x})\end{split}$ (13)
@@
$\begin{split}h_{i}(\bm{x})=\gamma^{i+1}\mathbb{E}_{\pi}[\prod_{j=1}^{i}1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(j))v(\bm{\phi}_{\pi}^{\bm{x}}(i+1))]\end{split}.$
Since $v(\cdot):\mathbb{R}^{n}\rightarrow\mathbb{R}$ is bounded over $\mathbb{R}^{n}$, we have $\lim_{i\rightarrow\infty}h_{i}(\bm{x})=0$ for $\bm{x}\in\mathbb{R}^{n}$ and consequently, $v(\bm{x})=\tilde{V}_{\gamma}(\bm{x})$ over $\mathbb{R}^{n}$. ∎
We can construct a necessary and sufficient barrier-like condition for the reach-avoid verification in Definition 4 via relaxing the Bellman equation (12), under the assumption that the reach-avoid probability $\mathbb{P}_{\pi}(RA_{\bm{x}_{0}})$ is strictly larger than the threshold $\epsilon_{2}$. This condition is the stochastic version of the one in Corollary 1 in [30].
@@
The reach-avoid probability $\mathbb{P}_{\pi}(RA_{\bm{x}_{0}})$ is strictly larger than the threshold $\epsilon_{2}$, i.e., $\mathbb{P}_{\pi}(RA_{\bm{x}_{0}})>\epsilon_{2}$, where $\epsilon_{2}\in[0,1)$.

### v3

@@ IV-A Reach-avoid Verification I
$\begin{split}\zeta_{k}(\bm{x}):&=\underbrace{\mathbb{E}_{\pi}\left[\sum_{i=0}^{k}\prod_{j=0}^{i-1}1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(j))\cdot 1_{\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(i))\right]}_{V_{k}(\bm{x})}+\mathbb{E}_{\pi}\left[\prod_{j=0}^{k}1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(j))\cdot v(\bm{\phi}_{\pi}^{\bm{x}}(k+1))\right]=v(\bm{x}).\end{split}$
Base case $k=0$:
$\begin{split}\zeta_{0}(\bm{x})&=\mathbb{E}_{\pi}\left[1_{\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(0))\right]+\mathbb{E}_{\pi}\left[1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(0))\cdot v(\bm{\phi}_{\pi}^{\bm{x}}(1))\right]\\ &=1_{\mathcal{X}_{r}}(\bm{x})+1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{x})\cdot\mathbb{E}_{\bm{\theta}}\left[v(\bm{f}(\bm{x},\bm{\theta}))\right]=v(\bm{x}),\end{split}$
where the last equality follows from the fixed-point condition on $v$.
Inductive step: Assume the statement holds for some $k\geq 0$, i.e., $\zeta_{k}(\bm{x})=v(\bm{x})$. Then,
$\zeta_{k+1}(\bm{x})=$ $\mathbb{E}_{\pi}\left[\sum_{i=0}^{k+1}\prod_{j=0}^{i-1}1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(j))\cdot 1_{\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(i))\right]+\mathbb{E}_{\pi}\left[\prod_{j=0}^{k+1}1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(j))\cdot v(\bm{\phi}_{\pi}^{\bm{x}}(k+2))\right]$
$=$ $\underbrace{\mathbb{E}_{\pi}\left[\sum_{i=0}^{k}\prod_{j=0}^{i-1}1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(j))\cdot 1_{\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(i))\right]}_{=V_{k}(\bm{x})}+\mathbb{E}_{\pi}\left[\prod_{j=0}^{k}1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(j))\cdot 1_{\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(k+1))\right]$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad+\mathbb{E}_{\pi}\left[\prod_{j=0}^{k+1}1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(j))\cdot v(\bm{\phi}_{\pi}^{\bm{x}}(k+2))\right]$
$=$ $\zeta_{k}(\bm{x})-\mathbb{E}_{\pi}\left[\prod_{j=0}^{k}1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(j))\cdot v(\bm{\phi}_{\pi}^{\bm{x}}(k+1))\right]$
$\quad+\mathbb{E}_{\pi}\Bigg[\prod_{j=0}^{k}1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(j))\cdot\Big(1_{\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(k+1))+1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(k+1))\cdot\mathbb{E}_{\bm{\theta}}\left[v(\bm{\phi}_{\pi}^{\bm{x}}(k+2))\right]\Big)\Bigg]$
$=$ $\zeta_{k}(\bm{x})-1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{x})h_{k}(\bm{x})+1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{x})h_{k}(\bm{x})$
$=$ $\zeta_{k}(\bm{x})=v(\bm{x}),$
where $v(\bm{\phi}_{\pi}^{\bm{x}}(k+1))=1_{\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(k+1))+1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(k+1))\cdot\mathbb{E}_{\bm{\theta}}\left[v(\bm{\phi}_{\pi}^{\bm{x}}(k+2))\right]$, which can be obtained via the Bellman equation (7).
By induction, $\zeta_{k}(\bm{x})=v(\bm{x})$ for all $k\geq 0$.
Finally, taking the limit as $k\to\infty$, we have
$\begin{split}v(\bm{x})&=\lim_{k\to\infty}\zeta_{k}(\bm{x})\\ &=\underbrace{\mathbb{E}_{\pi}\left[\sum_{i=0}^{\infty}\prod_{j=0}^{i-1}1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(j))\cdot 1_{\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(i))\right]}_{=V(\bm{x})}+1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{x})\lim_{k\to\infty}h_{k}(\bm{x}),\end{split}$
which establishes
$v(\bm{x})=V(\bm{x})+1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{x})\lim_{k\to\infty}h_{k}(\bm{x}).$ (9)
In the following, we show $\lim_{k\rightarrow\infty}h_{k}(\bm{x})=0$ for $\bm{x}\in\mathcal{X}\setminus\mathcal{X}_{r}$.
1). For all $\bm{x}\in\mathcal{X}\setminus\mathcal{X}_{r}$, the system (1) exits $\mathcal{X}\setminus\mathcal{X}_{r}$ in finite time almost surely, i.e.,
$\mathbb{P}_{\pi}\left(\forall k\in\mathbb{N}.\bm{\phi}^{\bm{x}}_{\pi}(k)\in\mathcal{X}\setminus\mathcal{X}_{r}\right)=0.$
This implies that trajectories starting in $\mathcal{X}\setminus\mathcal{X}_{r}$ will almost surely either
  1. 1.
    enter the target set $\mathcal{X}_{r}$, or
  2. 2.
    leave the safe set $\mathcal{X}$ (i.e., enter $\mathbb{R}^{n}\setminus\mathcal{X}$)
within finite time.
2). The function $v(\cdot)$ is bounded over $\mathbb{R}^{n}$. Thus, there exists a constant $M>0$ such that
$|v(\bm{y})|\leq M,\quad\forall\bm{y}\in\mathbb{R}^{n}.$
3). Define the event:
$A_{k}(\bm{x}):=\left\{\bm{\phi}^{\bm{x}}_{\pi}(j)\in\mathcal{X}\setminus\mathcal{X}_{r}\text{ for all }j=1,2,\dots,k\right\}.$
This represents the set of disturbance signals $\pi$ where the trajectory remains in $\mathcal{X}\setminus\mathcal{X}_{r}$ from time $1$ to $k$.
4). Assumption 1 implies that the event $\bigcap_{i=1}^{\infty}A_{i}(\bm{x})$ has probability zero, i.e.,
$\mathbb{P}_{\pi}\left(\bigcap_{k=1}^{\infty}A_{k}(\bm{x})\right)=\mathbb{P}_{\pi}\left(\forall k\geq 1.\bm{\phi}^{\bm{x}}_{\pi}(k)\in\mathcal{X}\setminus\mathcal{X}_{r}\right)=0.$
Since $A_{k+1}(\bm{x})\subseteq A_{k}(\bm{x})$ (the sequence is nested), continuity of probability measures gives
$\lim_{k\to\infty}\mathbb{P}_{\pi}\left(A_{k}(\bm{x})\right)=\mathbb{P}_{\pi}\left(\bigcap_{k=1}^{\infty}A_{k}(\bm{x})\right)=0.$
5). The term $\prod_{j=1}^{k}1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{\phi}^{\bm{x}}_{\pi}(j))$ is the indicator of $A_{k}(\bm{x})$. Using the boundedness of $v$ (assume $|v|\leq M$ over $\mathcal{X}\setminus\mathcal{X}_{r}$), we have
$\begin{split}|h_{k}(\bm{x})|&\leq\mathbb{E}_{\pi}\left[\prod_{j=1}^{k}1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(j))\cdot|v(\bm{\phi}_{\pi}^{\bm{x}}(k+1)|\right]\leq M\cdot\mathbb{E}_{\pi}\left[\prod_{j=1}^{k}1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(j))\right]=M\cdot\mathbb{P}_{\pi}\left(A_{k}(\bm{x})\right).\end{split}$
As $k\to\infty$, we have
$\lim_{k\to\infty}|h_{k}(\bm{x})|\leq M\cdot\lim_{k\to\infty}\mathbb{P}_{\pi}\left(A_{k}(\bm{x})\right)=0.$
Hence, $\lim_{k\to\infty}h_{k}(\bm{x})=0$ for $\bm{x}\in\mathcal{X}\setminus\mathcal{X}_{r}$. Consequently, $h_{k}(\bm{x})=0$ for $\bm{x}\in\mathcal{X}\setminus\mathcal{X}_{r}$.
Finally, when $\bm{x}\in\mathcal{X}_{r}$ or $\bm{x}\in\mathbb{R}^{n}\setminus\mathcal{X}$, the term $1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{x})$ in (9) is zero, thus $\lim_{k\rightarrow\infty}1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{x})h_{k}(\bm{x})=0$.
Consequently, $v(\bm{x})=V(\bm{x})$ over $\mathbb{R}^{n}$. ∎
@@
Since $v(\bm{x})$ satisfies (10), by following the inductive argument used in the proof of Proposition 4-where we showed
$v(\bm{x})=V(\bm{x})+1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{x})\lim_{k\to\infty}h_{k}(\bm{x})$
-but replacing the equality “=” with $\leq$”, we obtain that
$\begin{split}v(\bm{x})\leq V(\bm{x})+1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{x})\lim_{k\rightarrow\infty}h_{k}(\bm{x})\end{split}$
for $\bm{x}\in\mathbb{R}^{n}$, where $V(\cdot):\mathbb{R}^{n}\rightarrow\mathbb{R}$ is the value function defined in (6). Since $v(\bm{\phi}_{\pi}^{\bm{x}}(k+1))\leq 0$ when $\bm{\phi}_{\pi}^{\bm{x}}(k+1)\in\mathbb{R}^{n}\setminus\mathcal{X}$, we have
$\begin{split}h_{k}(\bm{x})&=\mathbb{E}_{\pi}[\prod_{j=1}^{k}1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(j))v(\bm{\phi}_{\pi}^{\bm{x}}(k+1))]\leq\mathbb{E}_{\pi}[\prod_{j=1}^{k}1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(j))w_{k+1}(\bm{x})],\end{split}$
where $w_{k+1}(\bm{x})=1_{\mathcal{X}}(\bm{\phi}_{\pi}^{\bm{x}}(k+1))v(\bm{\phi}_{\pi}^{\bm{x}}(k+1))$. Also, since $v(\cdot):\mathbb{R}^{n}\rightarrow\mathbb{R}$ is bounded over $\mathcal{X}$ and $\mathbb{P}_{\pi}(\forall k\in\mathbb{N}.\bm{\phi}_{\pi}^{\bm{x}}(k)\in\mathcal{X}\setminus\mathcal{X}_{r})=0$ for $\bm{x}\in\mathcal{X}\setminus\mathcal{X}_{r}$, we conclude $\lim_{k\rightarrow\infty}1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{x})h_{k}(\bm{x})=0$ for $\bm{x}\in\mathbb{R}^{n}$. Consequently, $v(\bm{x})\leq V(\bm{x})$ for $\bm{x}\in\mathbb{R}^{n}$.
@@
Remark 3
There is an important distinction between Proposition 3 and Theorem 2 that we now clarify explicitly here:
1) Proposition 3 provides a condition for verifying upper bounds on reach-avoid probabilities and does not require Assumption 1. This makes it broadly applicable, particularly in settings where the system may remain within $\mathcal{X}\setminus\mathcal{X}_{r}$ indefinitely with nonzero probability.
2) Theorem 2, on the other hand, provides necessary and sufficient conditions for verifying lower bounds on reach-avoid probabilities. However, it relies on Assumption 1, which ensures that the probability of the system (1) staying in $\mathcal{X}\setminus\mathcal{X}_{r}$ for all time is zero. This assumption is essential to guarantee that, if there exists a function satisfying condition (10), then the specified threshold $\epsilon_{2}$ is indeed a valid lower bound for the reach-avoid probability $\mathbb{P}_{\pi}(RA_{\bm{x}_{0}})$. Without Assumption 1, we cannot use condition (10) to justify lower bounds in the reach-avoid verification, since we cannot guarantee $\lim_{i\rightarrow\infty}h_{i}(\bm{x})=0$ for $\bm{x}\in\mathcal{X}\setminus\mathcal{X}_{r}$.
@@ IV-B Reach-avoid Verification II
The subsection will formulate a necessary and sufficient barrier-like condition for the reach-avoid verification without Assumption 1. Instead, another assumption that $\epsilon_{2}$ is strictly smaller than the exact reach-avoid probability $\mathbb{P}_{\pi}(RA_{\bm{x}_{0}})$ is imposed. Similar to the one in Subsection IV-A, this condition is constructed by relaxing a Bellman equation, which is derived from a discounted value function.
@@
$\begin{split}&\sum_{k=m+1}^{M}\mathbb{E}_{\pi}[\gamma^{k}\prod_{j=0}^{k-1}1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(j))1_{\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(k))]\leq\sum_{k=m+1}^{M}\mathbb{E}_{\pi}[\prod_{j=0}^{k-1}1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(j))1_{\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(k))]\end{split}$
@@
We first show that
$v(\bm{x})=\tilde{V}_{\gamma}(\bm{x})+1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{x})\lim_{k\to\infty}h_{k}(\bm{x})$
@@
$h_{k}(\bm{x}):=\gamma^{k+1}\mathbb{E}_{\pi}\left[\prod_{j=1}^{k}1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(j))\cdot v(\bm{\phi}_{\pi}^{\bm{x}}(k+1))\right].$
We note that when $k=0$, the product is taken over an empty index set, and by convention, the empty product equals 1. Therefore, $h_{0}(\bm{x})=\gamma\mathbb{E}_{\pi}[v(\bm{\phi}_{\pi}^{\bm{x}}(1))]=\gamma\mathbb{E}_{\bm{\theta}}[v(\bm{f}(\bm{x},\bm{\theta}))].$
For this sake, we prove by induction that for all $k\in\mathbb{N}$,
$\begin{split}\zeta_{k}(\bm{x}):&=\underbrace{\mathbb{E}_{\pi}\left[\sum_{i=0}^{k}\gamma^{i}\prod_{j=0}^{i-1}1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(j))\cdot 1_{\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(i))\right]}_{=\tilde{V}_{k,\gamma}(\bm{x})}+\gamma^{k+1}\mathbb{E}_{\pi}\left[\prod_{j=0}^{k}1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(j))\cdot v(\bm{\phi}_{\pi}^{\bm{x}}(k+1))\right]\\ &=v(\bm{x}).\end{split}$
Base case $k=0$:
$\begin{split}\zeta_{0}(\bm{x})&=\mathbb{E}_{\pi}\left[1_{\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(0))\right]+\gamma\mathbb{E}_{\pi}\left[1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(0))\cdot v(\bm{\phi}_{\pi}^{\bm{x}}(1))\right]\\ &=1_{\mathcal{X}_{r}}(\bm{x})+\gamma 1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{x})\cdot\mathbb{E}_{\bm{\theta}}\left[v(\bm{f}(\bm{x},\bm{\theta}))\right]\\ &=v(\bm{x}),\end{split}$
where the last equality follows from the Bellman equation (12).
Inductive step: Assume the statement holds for some $k\geq 0$, i.e., $\zeta_{k}(\bm{x})=v(\bm{x})$. Then,
$\zeta_{k+1}(\bm{x})$
$=\mathbb{E}_{\pi}\left[\sum_{i=0}^{k+1}\gamma^{i}\prod_{j=0}^{i-1}1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(j))\cdot 1_{\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(i))\right]+\gamma^{k+2}\mathbb{E}_{\pi}\left[\prod_{j=0}^{k+1}1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(j))\cdot v(\bm{\phi}_{\pi}^{\bm{x}}(k+2))\right]$
$=\underbrace{\mathbb{E}_{\pi}\left[\sum_{i=0}^{k}\gamma^{i}\prod_{j=0}^{i-1}1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(j))\cdot 1_{\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(i))\right]}_{=\tilde{V}_{k,\gamma}(\bm{x})}+\mathbb{E}_{\pi}\left[\gamma^{k+1}\prod_{j=0}^{k}1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(j))\cdot 1_{\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(k+1))\right]$
$\quad+\gamma^{k+2}\mathbb{E}_{\pi}\left[\prod_{j=0}^{k+1}1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(j))\cdot v(\bm{\phi}_{\pi}^{\bm{x}}(k+2))\right]$
$=\zeta_{k}(\bm{x})-\gamma^{k+1}\mathbb{E}_{\pi}\left[\prod_{j=0}^{k}1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(j))\cdot v(\bm{\phi}_{\pi}^{\bm{x}}(k+1))\right]$
$\quad\quad+\gamma^{k+1}\mathbb{E}_{\pi}\Bigg[\prod_{j=0}^{k}1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(j))\cdot\Big(1_{\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(k+1))+\gamma 1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(k+1))\cdot\mathbb{E}_{\bm{\theta}}\left[v(\bm{\phi}_{\pi}^{\bm{x}}(k+2))\right]\Big)\Bigg]$
$=\zeta_{k}(\bm{x})-1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{x})h_{k}(\bm{x})+1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{x})h_{k}(\bm{x})$
$=\zeta_{k}(\bm{x})=v(\bm{x}).$
By induction, $\zeta_{k}(\bm{x})=v(\bm{x})$ for all $k$.
Finally, taking the limit as $k\to\infty$, we have
$\begin{split}v(\bm{x})=\lim_{k\to\infty}\zeta_{k}(\bm{x})=\underbrace{\mathbb{E}_{\pi}\left[\sum_{i=0}^{\infty}\gamma^{i}\prod_{j=0}^{i-1}1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(j))\cdot 1_{\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(i))\right]}_{=\tilde{V}_{\gamma}(\bm{x})}+1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{x})\lim_{k\to\infty}h_{k}(\bm{x}),\end{split}$
which establishes
$v(\bm{x})=\tilde{V}_{\gamma}(\bm{x})+1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{x})\lim_{k\to\infty}h_{k}(\bm{x}).$
Since $v(\cdot):\mathbb{R}^{n}\rightarrow\mathbb{R}$ is bounded over $\mathbb{R}^{n}$, we have $\lim_{k\rightarrow\infty}h_{k}(\bm{x})=0$ for $\bm{x}\in\mathbb{R}^{n}$ and consequently, $v(\bm{x})=\tilde{V}_{\gamma}(\bm{x})$ over $\mathbb{R}^{n}$. ∎
We can construct a necessary and sufficient barrier-like condition for the reach-avoid verification in Definition 4 by relaxing the Bellman equation (12), under the assumption that the reach-avoid probability $\mathbb{P}_{\pi}(RA_{\bm{x}_{0}})$ is strictly larger than the threshold $\epsilon_{2}$. This condition is the stochastic version of the one in Corollary 1 in [zhao2022inner].
@@
The reach-avoid probability $\mathbb{P}_{\pi}(RA_{\bm{x}_{0}})$ is strictly larger than the threshold $\epsilon_{2}$, i.e.,
$\mathbb{P}_{\pi}(RA_{\bm{x}_{0}})>\epsilon_{2},$
where $\epsilon_{2}\in[0,1)$.
Assumption 2 is not overly restrictive and does not generally compromise the practical utility of the proposed method. In practice, since $\mathbb{P}_{\pi}(RA_{\bm{x}_{0}})$ is unknown, it is rare for the threshold $\epsilon_{2}$ set by engineers to exactly match $\mathbb{P}_{\pi}(RA_{\bm{x}_{0}})$. Therefore, either $\epsilon_{2}$ tends to be larger or smaller than $\mathbb{P}_{\pi}(RA_{\bm{x}_{0}})$, with both cases occurring frequently. When $\epsilon_{2}$ exceeds $\mathbb{P}_{\pi}(RA_{\bm{x}_{0}})$, certification is not possible, as the claim becomes infeasible. In contrast, the case where $\epsilon_{2}$ is smaller than $\mathbb{P}_{\pi}(RA_{\bm{x}_{0}})$, which corresponds to Assumption 2, is the focus of this work.